OpenStudy (anonymous):

what is the slope of the line with the equation 2x+3y+6=0

OpenStudy (anonymous):

you have to put this in the y=mx+b format, can you do that?

OpenStudy (anonymous):

you can do that. amistre doesn't. in standard form \[ax+by+c=0\] slope is \[-\frac{a}{b}\]

OpenStudy (anonymous):

same thing of course, but your choice

OpenStudy (anonymous):

I prefer to memorize methods rather than formulas ;p

OpenStudy (anonymous):

All you need to do is rearrange this equation by adding and subtracting the terms so it looks similar to (I know this will look unprofessional but just think that # means "number") #y=#x+#. Then simplify things so it's just y=#x+#. You'll find that this equation is in the y=mx+b format. The m means the slope, and that's your answer! You just need to mess with the equation to find the answer.